OpenStudy (anonymous):

The chance of a day being rainy is 20%. What is the chance of there being two rainy days during a week?

OpenStudy (anonymous):

Start by finding out how many rainy day combos there are in a week. Then, take that number and multiply it by (%rainy)^2 and multiply it again by (%not rainy)^5.

OpenStudy (anonymous):

Wouldn't the rainy day combos be 7^7?

OpenStudy (anonymous):

There is an easy way to figure the combos. It's starts by finding the permutations first. I'll explain.

OpenStudy (anonymous):

Look at the days, Sunday thru Saturday, as days 1-7. You could have the any of the days 1-7 as the "first" rainy day, so you start with a factor of 7.

OpenStudy (anonymous):

Now, ther are six days left for the "second" day (which may or may not precede the "first" day, hence the apostrophe's).

OpenStudy (anonymous):

So, there are 6 days from which to choose from for the "second", so the next step is 7 x 6, but you don't want to double count the pairs, so you have to divide by 2, so 7 x 6 / 2.

OpenStudy (anonymous):

7 x 6 / 2 is the number of combinations of 2 rainy day combos. There are a couple of other ways to get to this number. Perhaps the easiest is to start with day 1 and realize that there are 6 other days it can be paired with "going forward". Then take day 2 and realize that there are 5 to be paired with, etc. 6 + 5 + 4 + 3 + 2 + 1 = 42

OpenStudy (anonymous):

So 21 days that have 2 rainy days?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Not 21 days have rainy days, but more accurately, there are 21 pairs of days that are rainy.

OpenStudy (anonymous):

But how many total combinations are there

OpenStudy (anonymous):

7! or 5040?

OpenStudy (anonymous):

That's step1. Now, the second part is covered in the first reply. Multiplying by the probability of rainys x non-rainys

OpenStudy (anonymous):

What you are really asking is how many total pairs of days are there in 7 objects (days, dice, balls, hoops, etc.). Just 21 ways to take 2 differing pairs of things from a set of 7 different things.

OpenStudy (anonymous):

So about .275

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

K, Thanks

OpenStudy (anonymous):

You got it right and very quickly I might add

OpenStudy (anonymous):

Would the process be similar if the question was find the probability of at least 2 rainy days?

OpenStudy (anonymous):

Then, you would have to add up 2 rainy days + 3 rainy days + 4 +... Easier would to be finding the probability of no rainy days + 1 rainy day and subtracting that from 1.

OpenStudy (anonymous):

the probability for P(rain on 0 days) = .209 and the the probability for P(rain on 1 day) = .367 so the answer would be 1-.209-.367 which equals .424, right?

OpenStudy (anonymous):

Yes, that's right. You're actually pretty good at probabilty. I don't think you are going to have much trouble with this subject.

OpenStudy (anonymous):

I say that because you got both of these answers very quickly with minimal help.

OpenStudy (anonymous):

Mind answering another?

OpenStudy (anonymous):

Ok, but I don't think I'm as fast as you are!

OpenStudy (anonymous):

Here goes: First flip a coin. If you got heads, roll a 6 sided dice. If you got tails, roll an 8 sided dice. What is the chance that you flipped heads if you rolled a 6?

OpenStudy (anonymous):

I can do this one, but you'll have to bear with my step-by-step.

OpenStudy (anonymous):

K, go for

OpenStudy (anonymous):

Start with a "backwards to forwards" mindset on this. That is what you have to do with what is essentially an example of Bayes' Theorem. (Uses conditional probability). If you roll a 6-sided die, the probability is 1/6 or 4/24. If you roll an 8-sided die, the prob is 1/8 or 3/24. The total probability is 7/24. So, 3/24 / 7/24 = 3/7. Now let me think about this to see if my logic was right.

OpenStudy (anonymous):

Got my 6 and 8 switched. 4/24 / 7/24 = 4/7.

OpenStudy (anonymous):

I could draw you a nice matrix for this. How do I use a drawing tool here?

OpenStudy (anonymous):

so would the answer br 4/7 or .571?

OpenStudy (anonymous):

I just saw the drawing tool button

OpenStudy (anonymous):

Yep, it's right. Thanks so much

OpenStudy (anonymous):

Yes, that's the answer, but you might want to spend some time with this and the drawing as it is essentially a harder problem than the first. I'm going to try to draw a matrix. It's my first one, so it will be sloppy.

OpenStudy (anonymous):

Would you mind answering another question at about 7:00 PST, I have a piano lesson now.

OpenStudy (anonymous):

If i'm around, just look for me.

OpenStudy (anonymous):

I already became a fan, so that should be easy.

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

good luck, Chopin.

OpenStudy (anonymous):

 So it is really (4/48) / (7/48) = 4/7 which is the same answer of 0.571. With this drawing, I just divided the numerator and denominator by 1/2 to clarify the true probabilities of getting a "6". Same answer, just more rigorous.